OpenStudy (anonymous):

How to solve this equation with homogeneous coefficients x(x^2+y^2)(ydx-xdy)+y^6dy=0

OpenStudy (anonymous):

its a differential equation problem

OpenStudy (lgbasallote):

i think you should be a little more specific...

OpenStudy (anonymous):

how to solve this using the theorem 1:if M (x,y) and N(x,y) are both homogeneous & of the same degree, the function M(x,y)/N(x,y) is homogeneous of degree zero

OpenStudy (lgbasallote):

ohh that

OpenStudy (anonymous):

& using also theorem 2 of homogeneous function theorem 2: If f(x,y0 is homogeneous of degree zero in x and y, f(x,y) is a function of y/x alone.

OpenStudy (lgbasallote):

have you tried starting this?

OpenStudy (anonymous):

i'm confused in what to use if its y=vx or x=vy & when to substitute it..

OpenStudy (anonymous):

i have here the steps in solving

OpenStudy (lgbasallote):

you have to change it to look like Mdx + Ndy = 0 first

OpenStudy (anonymous):

1. replace y by y=vx or x=vy 2. Factor out x or y to obtain DE in which the coefficients are functions of v alone. 3. use variable separable method. If the DE is not exact

OpenStudy (lgbasallote):

change. it. to. Mdx + Ndy = 0. first.

OpenStudy (anonymous):

how?

OpenStudy (lgbasallote):

expand x(x^2 + y^2)(ydx - xdy)

OpenStudy (lgbasallote):

tag me when you got the exapnsion :D

OpenStudy (anonymous):

@lgbasallote (x^3+xy+xy^2)(xydx-x^2dy)+y^6dy=0

OpenStudy (lgbasallote):

uhh 1) how did you get that 2) that's still not in the form Mdx + Ndy = 0 `

OpenStudy (anonymous):

i distributed x.

OpenStudy (anonymous):

@lgbasallote i really don't know how to start this can u help me? pls. can you show me me how to solve this?

OpenStudy (lgbasallote):

well..this is still algebra...distribution i mean... distribut x into (x^2 + y^2) first. that gives you (x^3 +xy^2) now FOIL it with (ydx - xdy) you'll get x^3 ydx + xy^3 dx - x^4 dy - x^2 y^2 dy so the whole equation would be \[x^3 y dx + xy^3 dx - x^4 dy - x^2 y^2 dy + y^6 dy = 0\] factor out dx and dy \[\implies (x^3y + xy^3)dx - (x^4 + x^2y^2 + y^6)dy = 0\] that is now in the form Mdx + Ndy = 0

OpenStudy (lgbasallote):

now do your homo thingy